# Nonstandard Universes

### v1

@@ A.1 First-Order Languages, Models and Satisfaction
A first-order language $\mathcal{L}=(\mathcal{R},\mathcal{F})$ consists of a collection of relation symbols (predicate symbols) $\mathcal{R}$ and function symbols $\mathcal{F}$. If =$\mathcal{R}=\{R_{1},\dots,R_{n}\}$ and =$\mathcal{F}=\{F_{1},\dots,F_{m}\}$ are finite sets, we may write $\mathcal{L}=(R_{1},\dots,R_{n},F_{1},\dots,F_{m})$.
@@
[…]
@@
If =$\mathcal{R}=\{R_{1},\dots,R_{n}\}$ and =$\mathcal{F}=\{F_{1},\dots,F_{m}\}$ are finite sets, we may write $\mathfrak{A}=(A,R_{1}^{\mathfrak{A}},\dots R_{n}^{\mathfrak{A}},F_{1}^{\mathfrak{A}},\dots,F_{m}^{\mathfrak{A}})$. When the interpretation is clear, we may dispense with the $\mathfrak{A}$-superscripts entirely, and simply write $(A,R_{1},\dots,R_{n},F_{1},\dots,F_{m})$.
@@
  • The equality symbol $=$, which is always to be interpreted as equality.
@@
  • $(\varphi\lor\psi)$ is an abbreviation for ¬($\lnot(\lnot\varphi\land\lnot\psi)$.
  • $(\varphi\to\psi)$ is an abbreviation for ($(\lnot\varphi\lor\psi)$.
@@
[…]
@@
      If $\varphi$ is the formula $\psi\land\chi$, then $\mathfrak{A}\vDash\varphi[a_{1},\dots,a_{n}]$ if and only if both $\mathfrak{A}\vDash\psi[a_{1},\dots,a_{n}]$ and $\mathfrak{A}\vDash\chi[a_{1},\dots,a_{n}]$
@@
for every $\Leftrightarrow\quad\text{for every $a\in A$, }\mathfrak{A}\vDash\psi[a_{1},\dots,a_{i-1},a,a_{i+1},\dots,a_{p}],$
[…]
@@ A.2 Elementary Embeddings and Elementary Equivalence
[…]
@@
[…]
@@
[…]
@@ A.3 Ultrafilters
    A family $\mathcal{A}\subseteq\mathcal{P}(I)$ is said to satisfy the finite intersection property (f.i.p.) if and only whenever :=$\mathcal{B}:=\{A_{1},\dots,A_{n}\}$ is any finite subcollection of $\mathcal{A}$, then $\bigcap\mathcal{B}$ is non-empty, i.e. $A_{1}\cap\dots\cap A_{n}\neq\varnothing$.
@@
In Section 6.1 it will be shown that any ultrafilter will induce a transfer map $*:V(X)\to V(Y)$ between superstructures. In order for that transfer map to give rise to a nonstandard framework — i.e. so that ${}^{\sigma}C\subsetneq{}^{*}C$ for some countable $C\in V(X)$ — we will require that the ultrafilter is countably incomplete.
@@
    :=$\mathcal{F}:=\{F\subseteq I:\text{there are }A_{1},\dots,A_{n}\in\mathcal{A}\text{ such that }A_{1}\cap\dots\cap A_{n}\subseteq F\}$
@@ A.4 Ultraproducts and Ultrapowers
[…]
@@ Appendix B Existence of Good Ultrafilters
Proof: The ()($\Rightarrow$)( ⇒ )-direction is obvious.
For the ()($\Leftarrow$)( ⇐ )-direction, suppose that $\alpha\leq\kappa$ and that $p:\mathcal{P}^{<\omega}(\alpha)\to\mathcal{U}$ is an order-reversal. Note that if $s$ is a finite subset of $\kappa$, then $s\cap\alpha$ is a finite subset of $\alpha$. Thus we can define an order-reversing map
@@
Observe that each $f_{J}:\kappa\to\kappa$ is surjective, as can be seen from ()($\star$)( ⋆ ) with $n=1$. Thus if we define $P_{J}$ to be the partition of $\kappa$ given by
@@
Let $\Pi:=\{P_{J}:J\subseteq\kappa\}$. To show that $(\Pi,\mathcal{E})$ is consistent, suppose that $E\in\mathcal{E}$ and that $C_{1},\dots,C_{n}$ belong to $P_{J_{1}},\dots,P_{J_{n}}$, where the $J_{i}$ are distinct subsets of $\kappa$. Then there exists $\alpha<\kappa$ such that $E=E_{\alpha}$, and there exist $\gamma_{1},\dots,\gamma_{n}<\kappa$ so that $C_{i}=f_{J_{i}}^{-1}\{\gamma_{i}\}$ for $i=1,\dots,n$. But then by ()($\star$)( ⋆ ) there exists $\xi\in I_{\alpha}$ such that $f_{J_{i}}(\xi)=\gamma_{i}$, from which it follows that
@@
:=$\mathcal{E}^{\prime}:=\{E\cap q(s):E\in\mathcal{E},s\in\mathcal{P}^{<\omega}(\kappa)\},$

### v2

@@ A.1 First-Order Languages, Models and Satisfaction
A first-order language $\mathcal{L}=(\mathcal{R},\mathcal{F})$ consists of a collection of relation symbols (predicate symbols) $\mathcal{R}$ and function symbols $\mathcal{F}$. If $\mathcal{R}=\{R_{1},\dots,R_{n}\}$ and $\mathcal{F}=\{F_{1},\dots,F_{m}\}$ are finite sets, we may write $\mathcal{L}=(R_{1},\dots,R_{n},F_{1},\dots,F_{m})$.
@@
($\mathcal{L}$-structure) Let $\mathcal{L}=(\mathcal{R},\mathcal{F})$ be a first-order language. An $\mathcal{L}$-structure is a tuple $\mathfrak{A}=(A,\mathcal{L}^{\mathfrak{A}})$ where $\mathcal{L}^{\mathfrak{A}}=(\mathcal{R}^{\mathfrak{A}},\mathcal{F}^{\mathfrak{A}})$ consists of relations and functions on the set $A$. Specifically, for each $n$-ary relation symbol $R\in\mathcal{R}$ there corresponds an $n$-ary relation $R^{\mathfrak{A}}\in\mathcal{R}^{\mathfrak{A}}$ on $A$, and to each $n$-ary function symbol $F\in\mathcal{F}$ there corresponds an $n$-ary function $F^{\mathfrak{A}}\in\mathcal{F}^{\mathfrak{A}}$ on $A$ such that
@@
If $\mathcal{R}=\{R_{1},\dots,R_{n}\}$ and $\mathcal{F}=\{F_{1},\dots,F_{m}\}$ are finite sets, we may write $\mathfrak{A}=(A,R_{1}^{\mathfrak{A}},\dots R_{n}^{\mathfrak{A}},F_{1}^{\mathfrak{A}},\dots,F_{m}^{\mathfrak{A}})$. When the interpretation is clear, we may dispense with the $\mathfrak{A}$-superscripts entirely, and simply write $(A,R_{1},\dots,R_{n},F_{1},\dots,F_{m})$.
@@
  • The equality symbol ==, which is always to be interpreted as equality.
@@
  • $(\varphi\lor\psi)$ is an abbreviation for $\lnot(\lnot\varphi\land\lnot\psi)$.
  • $(\varphi\to\psi)$ is an abbreviation for $(\lnot\varphi\lor\psi)$.
@@
(Interpretation and Satisfaction) Let $\mathfrak{A}=(A,\mathcal{L}^{\mathfrak{A}})$ be a model for a first-order language $\mathcal{L}$. For this definition, we will write $t(x_{1},\dots,x_{n})$ if $t$ is a term all of whose variables are among $x_{1},\dots,x_{n}$ — they need not all occur, however. Similarly, if all the variables , we write $\varphi(x_{1},\dots,x_{n})$ if $\varphi$ is a formula all of whose variables are among $(x_{1},\dots,x_{n})$ — again, they need not all occur.
@@
      If $\varphi$ is the formula $\psi\land\chi$, then $\mathfrak{A}\vDash\varphi[a_{1},\dots,a_{n}]$ if and only if both $\mathfrak{A}\vDash\psi[a_{1},\dots,a_{n}]$ and $\mathfrak{A}\vDash\chi[a_{1},\dots,a_{n}]$ .
@@
$\Leftrightarrow\quad\text{for every $a\in A$, }\mathfrak{A}\vDash\psi[a_{1},\dots,a_{i-1},a,a_{i+1},\dots,a_{p}],$
$\Leftrightarrow\quad\text{for every $a\in A$, }\mathfrak{A}\vDash\psi[b_{1},\dots,b_{i-1},a,b_{i+1},\dots,b_{q}],\quad\text{(induction hypothesis)}$
@@ A.2 Elementary Embeddings and Elementary Equivalence
(Submodel) Suppose that $\mathfrak{A}=(A,\mathcal{L}^{\mathfrak{A}})$ and $\mathfrak{B}=(B,\mathcal{L}^{\mathfrak{B}})$ are models of a first-order language $\mathcal{L}$. We say that $\mathfrak{A}$ is a submodel of $\mathfrak{B}$ — and write $\mathfrak{A}\subseteq\mathfrak{B}$ — if and only if
@@
(Theory, Elementary Diagram) Suppose that $\mathfrak{A}=(A,\mathcal{L}^{\mathfrak{A}})$ is a model of a first-order language $\mathcal{L}$.
@@
Note that if $\varphi(x_{1},\dots,x_{n})$ is an $\mathcal{L}$–formula and $\mathfrak{A}=(A,\mathcal{L}^{\mathfrak{A}})$ is an $\mathcal{L}$-model, then for any $a_{1},\dots,a_{n}\in A$, we obtain an $\mathcal{L}_{A}$-sentence $\varphi(c_{a_{1}},\dots,c_{a_{n}})$ with the property that
@@ A.3 Ultrafilters
    A family $\mathcal{A}\subseteq\mathcal{P}(I)$ is said to satisfy the finite intersection property (f.i.p.) if and only whenever $\mathcal{B}:=\{A_{1},\dots,A_{n}\}$ is any finite subcollection of $\mathcal{A}$, then $\bigcap\mathcal{B}$ is non-empty, i.e. $A_{1}\cap\dots\cap A_{n}\neq\varnothing$.
@@
In Section 6.1 it will be shown that any ultrafilter will induce a transfer map $V(X)\to V(Y)$ between superstructures. In order for that transfer map to give rise to a nonstandard framework — i.e. so that $^{\sigma}C\subsetneq{}^{*}C$ for some countable $C\in V(X)$ — we will require that the ultrafilter is countably incomplete.
@@
    $\mathcal{F}:=\{F\subseteq I:\text{there are }A_{1},\dots,A_{n}\in\mathcal{A}\text{ such that }A_{1}\cap\dots\cap A_{n}\subseteq F\}$
@@ A.4 Ultraproducts and Ultrapowers
Suppose that $\mathfrak{A}=(A,\mathcal{L}^{\mathfrak{A}})$ is a model of a first-order language $\mathcal{L}$. Let $\mathcal{U}$ be an ultrafilter over the set $I$. For $a\in A$, let $c_{a}\in A^{I}$ be the constant map with value $a$.Then the map $h:\mathfrak{A}\to\mathfrak{A}^{I}/\mathcal{U}:a\mapsto c_{a}/\mathcal{U}$ is an elementary embedding.
@@ Appendix B Existence of Good Ultrafilters
Proof: The $(\Rightarrow)$-direction is obvious.
For the $(\Leftarrow)$-direction, suppose that $\alpha\leq\kappa$ and that $p:\mathcal{P}^{<\omega}(\alpha)\to\mathcal{U}$ is an order-reversal. Note that if $s$ is a finite subset of $\kappa$, then $s\cap\alpha$ is a finite subset of $\alpha$. Thus we can define an order-reversing map
@@
Observe that each $f_{J}:\kappa\to\kappa$ is surjective, as can be seen from $(\star)$ with $n=1$. Thus if we define $P_{J}$ to be the partition of $\kappa$ given by
@@
Let $\Pi:=\{P_{J}:J\subseteq\kappa\}$. To show that $(\Pi,\mathcal{E})$ is consistent, suppose that $E\in\mathcal{E}$ and that $C_{1},\dots,C_{n}$ belong to $P_{J_{1}},\dots,P_{J_{n}}$, where the $J_{i}$ are distinct subsets of $\kappa$. Then there exists $\alpha<\kappa$ such that $E=E_{\alpha}$, and there exist $\gamma_{1},\dots,\gamma_{n}<\kappa$ so that $C_{i}=f_{J_{i}}^{-1}\{\gamma_{i}\}$ for $i=1,\dots,n$. But then by $(\star)$ there exists $\xi\in I_{\alpha}$ such that $f_{J_{i}}(\xi)=\gamma_{i}$, from which it follows that
@@
$\mathcal{E}^{\prime}:=\{E\cap q(s):E\in\mathcal{E},s\in\mathcal{P}^{<\omega}(\kappa)\},$